A trough is 12 feet long, 3 feet deep, and 3 feet wide. Find the depth of the water when the trough contains 70 gallons(1 gallon is about 0.13368 cubic foot).
Answer by checkley77(12844)

12*3*x=36x ft^3 is the volume to be filled.
36x=70*.13368
36x=9.3576
x=9.3576/36
x=.2599 ft. is the depth of the 70 gallons of water.
Proof:
12*3*.2599=70*.13368
9.3564=9.3576